# RAISE: Realness Assessment for Image Synthesis and Evaluation

### v1

@@ I Introduction
The emergence of diffusion based generative models has revolutionized the creation of synthetic visual data, with unprecedented levels of photorealism being attained by models in this class such as Stable diffusion [1] and DALL-E 2  [2]. The ability of such generative models to process complex textual and visual prompts to produce photorealistic results distinguishes them from their predecessors, such as those based on generative adversarial learning or variational autoencoders. The widespread availability of such sophisticated models that can create visual data that are virtually indistinguisable from photographic depictions of real scenes raises concerns regarding privacy violations, intellectual property issues, security threats, and the spread of misinformation. To address such concerns, researchers have focused on developing robust solutions to distinguish AI generated images and videos from real ones.
@@
Despite such challenges, photorealistic synthetic data could be invaluable in numerous scenarios, where acquiring or accessing real data is difficult, expensive or intrusive. While large scale image and video datasets are available, public datasets of similar scales are quite rare for domains like biomedical imaging, as acquiring such images are often expensive as well as intrusive, while distributing them is likely to be precluded by protocols that are designed to protect the privacy of individual test subjects. Likewise, high-resolution satellite images, underwater images and astronomical images require highly specialized imaging equipment, making them costly to acquire. Further, even images and videos of natural scenes such as movie footages may be restricted by copyrights, limiting their use for most purposes. Such limitations on the amount of available data often stymies the development of robust deep learning models. Generative AI, capable of producing realistic and domain-specific visual data, offers a promising solution to this problem by augmenting the available data with generated samples. However, this strategy relies on the ability to quantify how closely synthetic data resemble real images. While humans can often identify subtle visual artifacts produced by AI, human evaluation is subjective and unscalable. Thus, developing reliable models to quantify perceptual realness is crucial for the practical deployment of AI generated visual content—a direction that remains underexplored.
@@
To address the challenge of quantifying the perceptual realness of AI generated visual content, we introduce RAISE, a dataset of AI generated images annotated with their subjective realness scores collected via a systematic human study. As humans are the primary consumers of visual media, this dataset provides a valuable resource for training and benchmarking models that estimate perceived realness. Such models can further guide the fine-tuning of generative AI models to reduce visual inconsistencies and produce more realistic images. In this paper, we decribe the construction of then RAISE dataset, analyze it statistically, and use it to build baseline models for realness prediction [^1] Our dataset and realness prediction models are avaialable at https://github.com/annimukherjee/RAISE.git.
@@ III Dataset Construction
In this paper, we introduce a dataset for Realness Assessment for Image Synthesis and Evaluation (RAISE), which comprises subjective realness scores for 600 images. Unlike existing datasets such as [8, 9, 11], which provide perceptual quality scores, RAISE focuses specifically on perceptual realness—a distinct visual concept. While perceptual quality is influenced by technical factors like distortions or visual artifacts, perceptual realness is dictated by the plausibility of the depicted scene, logical object placement, and natural appearance of colors and textures. Although quality degradations may affect perceived realness, they play a smaller role compared to semantic and contextual coherence, as shown in [10]. Indeed, camera captured images often contain distortions that lower technical quality without necessarily reducing perceived realness. To explore this relationship, RAISE also includes real images, allowing us to collect and compare realness ratings for both real and synthetic content. We argue that perceptual realness should be studied independently of technical quality, especially given that established IQA/VQA models for natural images can be readily applied to AIGIs to assess their technical quality when such AIGIs are realistic, i.e. they share similar distribution with real images. Thus, the purpose of the RAISE dataset is to facilitate the development and benchmarking of objective realness assessment models for AIGIs.
@@ III-A Content Collection
Both the real and AI generated images were carefully curated to encompass a broad range of scenes. Specifically, as human faces are highly photographed and salient visual entitites that are also particularly vulnerable to AI generated visual distortions, we ensured that the AIGI content of our dataset has a balanced representation of images depicting human faces. AIGIs were manually chosen to span a wide range of perceptual realness levels. The FID score calculated between the real and AIGI subsets of our dataset is 177.07. Both the real and the AIGI subsets of RAISE exhibit a wide range of scene complexities, as illustrated by the distribution of the spatial information (SI) [24] values in Fig. 1
@@ III-B Design of Subjective Study
We designed a psychovisual human study to collect reliable subjective ratings of the perceived realness following the ITU recommendation outlined in [25]. The human study was conducted using a single stimulus testing methodology, where each participant rated the perceived realness of a single image–either real or AIGI–on a continuos scale ranging from 0-100, labeled with absolute category rating (ACR) labels to describe the realness levels. The rating interface is demonstrated in Fig. 2.
@@
A total of 47 subjects participated in the study, comprising 41 men and 6 women, aged between 20 and 31 years (mean age: 23.5 years). Additional participant details that were obtained using standard tests for visual acuity and color vision as well as through a pre-study quiestionnaire about prior exposure to AI generated imagery are reported in Table I.
@@
The human study described in this section resulted in 14,100 total ratings, with each image’s realnes assessed by at least 23 distinct participants.
@@ V-A Feature Design
    Texture and Structure Features: AIGI are often characterized by unnatural textures such as unusual smoothness, repetitive patterns, and inconsistent spatial structures [28]. We derive texture descriptors from gray level cooccurrence matrix (GLCM) [29] and structure descriptors from histogram of oriented gradients (HoG) [30] to account for such effects [31].
@@
We trained various machine learning models using 55 handcrafted features across all the categories described above, computed from images in the test set of the RAISE dataset. Among the models evaluated, a decision tree regressor yielded the best performance. Table II reports the average performance of this regressor on the test partition of the RAISE dataset. The model achieved a SROCC of 0.5033 and a Pearson’s linear correlation coefficient (PLCC) of 0.4339, as shown in the last row of the table. These results demonstrate that even a compact set of relatively simple, low and mid-level features can capture perceptual cues influencing human judgments of image realness to a reasonable extent.
@@
Category # of Features Performance (excluding col. 1 for rows 1–5)
@@ V-C Transfer Learning
Transfer learning with foundation vision models is particularly effective when annotated data is limited, as in our case. For the purpose of transfer learning, we selected the ResNet-18 [35] backbone pretrained on ImageNet [36], which is well-suited for feature extraction across diverse vision tasks. We used the 512-dimensional feature map obtained by performing global average pooling of the deep features produced by the last residual block of ResNet-18, as the input to a neural network (NN) with 4 fully connected layers. This NN regressor conisted of 230,145 trainable parameters. We used dropout and data augmentation for regularization as described in V-B and the parameters of the ResNet-18 backbone remained unchanged during training.
@@ V-D JOINT
Since the JOINT model [10] was trained on subjective opinions of rationality and technical quality to predict overall AIGI naturalness, it is pertinent to evaluate its performance on the RAISE daataset introduced in this paper. We thus conducted inference on the RAISE test set with the trained JOINT model with official weights  [^2] Official JOINT implementation: https://github.com/zijianchen98/AGIN. Since the JOINT model predicts naturalness scores in the range of 1-5, we compared the model’s predictions with the subjective MOS ratings by scaling the MOS ratings to the same range.
@@
We next applied transfer learning using the pretrained JOINT model parameters. Since our dataset contains subjective realness annotations, we utlilized the rationality branch of JOINT which incorporates a fine-tuned ResNet-50 backbone pretrained on the AVA dataset [5]. We used the learned weights of JOINT to perform feature extraction with this backbone network, and trained a fully connected NN regressor having 32,582,501 parameters with dropout and data augmentation regularization. This transfer learning approach yielded an SROCC of 0.6785 and a PLCC of 0.7027 on the test set, as reported in the 4th row of Table III While the SROCC improved, the PLCC declined relative to transfer learning with ResNet-18 initialized with the default set of parameters trained on ImageNet [36]
We further fine-tuned the parameters of all the layers of the last residual blocks of the rationality branch of JOINT while training the fully connected NN regressor from scratch, resulting in a total of 47,547,237 parameters, traine with dropout and data augmentation regularization. The resulting performance is reported in the last row of Table III. The SROCC improved to 0.6798, while the PLCC improved to 0.7421. A scatter plot of the predicted versus ground-truth MOS scores obtained with this model is shown in Fig. 7. These results demonstrate that although the pretrained JOINT model performed poorly on the RAISE dataset, the features learned by its rationality branch significantly improved prediction performance when adapted through partial fine-tuning.
@@
Model # of Parameters Trained Performance
@@
NN on fine-tuned JOINT rationality features 47,547,237 0.6798 0.7421

### v2

@@ I Introduction
The emergence of diffusion based generative models has revolutionized the creation of synthetic visual data, with unprecedented levels of photorealism being attained by models in this class such as Stable diffusion [1] and DALL-E 2  [2]. The ability of such generative models to process complex textual and visual prompts to produce photorealistic results distinguishes them from their predecessors, such as those based on generative adversarial learning or variational autoencoders. The widespread availability of such sophisticated models that can create visual data that are virtually indistinguishable from photographic depictions of real scenes raises concerns regarding privacy violations, intellectual property issues, security threats, and the spread of misinformation. To address such concerns, researchers have focused on developing robust solutions to distinguish AI generated images and videos from real ones.
@@
Despite such challenges, photorealistic synthetic data could be invaluable in numerous scenarios, where acquiring or accessing real data is difficult, expensive or intrusive. While large scale image and video datasets are available, public datasets of similar scales are quite rare for domains like biomedical imaging, as acquiring such images are often expensive as well as intrusive, while distributing them is likely to be precluded by protocols that are designed to protect the privacy of individual test subjects. Likewise, high-resolution satellite images, underwater images and astronomical images require highly specialized imaging equipment, making them costly to acquire. Further, even images and videos of natural scenes such as movie footage may be restricted by copyrights, limiting their use for most purposes. Such limitations on the amount of available data often stymies the development of robust deep learning models. Generative AI, capable of producing realistic and domain-specific visual data, offers a promising solution to this problem by augmenting the available data with generated samples. However, this strategy relies on the ability to quantify how closely synthetic data resemble real images. While humans can often identify subtle visual artifacts produced by AI, human evaluation is subjective and unscalable. Thus, developing reliable models to quantify perceptual realness is crucial for the practical deployment of AI generated visual content—a direction that remains underexplored.
@@
To address the challenge of quantifying the perceptual realness of AI generated visual content, we introduce RAISE, a dataset of AI generated images annotated with their subjective realness scores collected via a systematic human study. As humans are the primary consumers of visual media, this dataset provides a valuable resource for training and benchmarking models that estimate perceived realness. Such models can further guide the fine-tuning of generative AI models to reduce visual inconsistencies and produce more realistic images. In this paper, we describe the construction of then RAISE dataset, analyze it statistically, and use it to build baseline models for realness prediction [^1] Our dataset and realness prediction models are available at https://github.com/annimukherjee/RAISE.git.
@@ III Dataset Construction
In this paper, we introduce a dataset for Realness Assessment for Image Synthesis and Evaluation (RAISE), which comprises subjective realness scores for 600 images. Unlike existing datasets such as [8, 9, 11], which provide perceptual quality scores, RAISE focuses specifically on perceptual realness—a distinct visual concept. While perceptual quality is influenced by technical factors like distortions or visual artifacts, perceptual realness is dictated by the plausibility of the depicted scene, logical object placement, and natural appearance of colors and textures. Although quality degradation may affect perceived realness, they play a smaller role compared to semantic and contextual coherence, as shown in [10]. Indeed, camera captured images often contain distortions that lower technical quality without necessarily reducing perceived realness. To explore this relationship, RAISE also includes real images, allowing us to collect and compare realness ratings for both real and synthetic content. We argue that perceptual realness should be studied independently of technical quality, especially given that established IQA/VQA models for natural images can be readily applied to AIGIs to assess their technical quality when such AIGIs are realistic, i.e. they share similar distribution with real images. Thus, the purpose of the RAISE dataset is to facilitate the development and benchmarking of objective realness assessment models for AIGIs.
@@ III-A Content Collection
Both the real and AI generated images were carefully curated to encompass a broad range of scenes. Specifically, as human faces are highly photographed and salient visual entities that are also particularly vulnerable to AI generated visual distortions, we ensured that the AIGI content of our dataset has a balanced representation of images depicting human faces. AIGIs were manually chosen to span a wide range of perceptual realness levels. The FID score calculated between the real and AIGI subsets of our dataset is 177.07. Both the real and the AIGI subsets of RAISE exhibit a wide range of scene complexities, as illustrated by the distribution of the spatial information (SI) [24] values in Fig. 1
@@ III-B Design of Subjective Study
We designed a psychovisual human study to collect reliable subjective ratings of the perceived realness following the ITU recommendation outlined in [25]. The human study was conducted using a single stimulus testing methodology, where each participant rated the perceived realness of a single image–either real or AIGI–on a continuous scale ranging from 0-100, labeled with absolute category rating (ACR) labels to describe the realness levels. The rating interface is demonstrated in Fig. 2.
@@
A total of 47 subjects participated in the study, comprising 41 men and 6 women, aged between 20 and 31 years (mean age: 23.5 years). Additional participant details that were obtained using standard tests for visual acuity and color vision as well as through a pre-study questionnaire about prior exposure to AI generated imagery are reported in Table I.
@@
The human study described in this section resulted in 14,100 total ratings, with each image’s realness assessed by at least 23 distinct participants.
@@ V-A Feature Design
    Texture and Structure Features: AIGI are often characterized by unnatural textures such as unusual smoothness, repetitive patterns, and inconsistent spatial structures [28]. We derive texture descriptors from gray level co-occurrence matrix (GLCM) [29] and structure descriptors from histogram of oriented gradients (HoG) [30] to account for such effects [31].
@@
We trained various machine learning models using 55 handcrafted features across all the categories described above, computed from images in the training partition of the RAISE dataset. Among the models evaluated, a decision tree regressor yielded the best performance. Table II reports the average performance of this regressor on the test partition of the RAISE dataset. The model achieved a SROCC of 0.5033 and a Pearson’s linear correlation coefficient (PLCC) of 0.4339, as shown in the last row of the table. These results demonstrate that even a compact set of relatively simple, low and mid-level features can capture perceptual cues influencing human judgments of image realness to a reasonable extent.
@@
Category
# of
Features
Performance
(excluding col. 1 for rows 1–5)
@@ V-C Transfer Learning
Transfer learning with foundation vision models is particularly effective when annotated data is limited, as in our case. For the purpose of transfer learning, we selected the ResNet-18 [35] backbone pretrained on ImageNet [36], which is well-suited for feature extraction across diverse vision tasks. We used the 512-dimensional feature map obtained by performing global average pooling of the deep features produced by the last residual block of ResNet-18, as the input to a neural network (NN) with 4 fully connected layers. This NN regressor consisted of 230,145 trainable parameters. We used dropout and data augmentation for regularization as described in V-B and the parameters of the ResNet-18 backbone remained unchanged during training.
@@ V-D JOINT
Since the JOINT model [10] was trained on subjective opinions of rationality and technical quality to predict overall AIGI naturalness, it is pertinent to evaluate its performance on the RAISE dataset introduced in this paper. We thus conducted inference on the RAISE test set with the trained JOINT model with official weights  [^2] Official JOINT implementation: https://github.com/zijianchen98/AGIN. Since the JOINT model predicts naturalness scores in the range of 1-5, we compared the model’s predictions with the subjective MOS ratings by scaling the MOS ratings to the same range.
@@
We next applied transfer learning using the pretrained JOINT model parameters. Since our dataset contains subjective realness annotations, we utilized the rationality branch of JOINT which incorporates a fine-tuned ResNet-50 backbone pretrained on the AVA dataset [5]. We used the learned weights of JOINT to perform feature extraction with this backbone network, and trained a fully connected NN regressor having 32,582,501 parameters with dropout and data augmentation regularization. This transfer learning approach yielded an SROCC of 0.6785 and a PLCC of 0.7027 on the test set, as reported in the 4th row of Table III While the SROCC improved, the PLCC declined relative to transfer learning with ResNet-18 initialized with the default set of parameters trained on ImageNet [36]
We further fine-tuned the parameters of all the layers of the last residual blocks of the rationality branch of JOINT while training the fully connected NN regressor from scratch, resulting in a total of 47,547,237 parameters, train with dropout and data augmentation regularization. The resulting performance is reported in the last row of Table III. The SROCC improved to 0.6798, while the PLCC improved to 0.7421. A scatter plot of the predicted versus ground-truth MOS scores obtained with this model is shown in Fig. 7. These results demonstrate that although the pretrained JOINT model performed poorly on the RAISE dataset, the features learned by its rationality branch significantly improved prediction performance when adapted through partial fine-tuning.
@@
Model
# of
Parameters
Trained
Performance
@@
NN on fine-tuned JOINT
rationality features
47,547,237 0.6798 0.7421